7.) Suppose that a fair coin is tossed 10 times and lands on heads exactly 2 times. Assuming that the tosses are independent, show that the conditional probability that the first toss landed on heads is 0.2. 8.) Suppose that X is uniformly distributed on [0,1] and let A be the event that X є 10,05) and let B be the event that X e [0.25,0.5) U[0.75,1.0). Show that A and B are independent.
Answer #1

(7) Let X be a r.v. of getting no. heads out of 10 tosses Then X~ Binomial(10,1/2) 10) 1 10 where, B be of event of getting 2

Required probability P BA) P(A) (10

8, Plan B) = P(0.25-X < 0.5) = /as dr = 0.25 P(A) = P(0 < X < 0.5) = dz 0.25 0.5 0 0.5 1 P(B) = P(0.25 < X < 0.5) + P(0.75 <